OpenStudy (anonymous):

Using r(t) = acos(t)i+asin(t)j, T(t) = (r' (t))/(||r' (t)||) and curvature is k(t) = (||T' (t) ||)/(||r' (t) ||) Prove that the curvature of a circle is constant and = 1/a

OpenStudy (jamesj):

So you just need to calculate. I.e., given r, what is r', ||r'||, T and k ? Just do the calculations; it's very straightforward.

OpenStudy (anonymous):

Really? Once I do the calculations you think i'll just see the proof?

OpenStudy (jamesj):

Just try it, really. Muster up your courage and do it.

OpenStudy (anonymous):

haha, ok, i shall return!

OpenStudy (jamesj):

Let's break this down. First, what's r' , ||r'|| and therefore T?

OpenStudy (jamesj):

I'm nervous you're typing for so long. Give it to me in pieces. What first is r' and ||r'||?

OpenStudy (jamesj):

even more nervous now.

OpenStudy (anonymous):

\[r(t)=acos(t)i+asin(t)j\] \[r'(t)=(a'\cos(t)-asin(t))i+(a'\sin(t)+acos(t))j\] \[||r'(t)||=\sqrt{(a'cost-asint)^{2}+(a'sint+acost)^{2}} \] \[=\sqrt{(a'\cos)^{2}-2a'acostsint+(asint)^{2}+(a'sint)^{2}+2a'acostsint+(acost)^{2}}\] \[=\sqrt{(a'\cos)^{2}+(asint)^{2}+(a'sint)^{2}+(acost)^{2}}\] \[=\sqrt{a'^{2}\cos^{2}t+a^{2}\sin^{2}t+a'^{2}\sin^{2}t+a^{2}\cos^{2}t}\]

OpenStudy (anonymous):

lol, oops just saw your messages

OpenStudy (anonymous):

sorry

OpenStudy (jamesj):

a is a constant.

OpenStudy (anonymous):

I believe T is called the tangent unit

OpenStudy (anonymous):

a is a consant, ok

OpenStudy (anonymous):

Am I on the right path though?

OpenStudy (anonymous):

ah ha!

OpenStudy (jamesj):

one sec on the phone. remember that a' =0 everywhere

OpenStudy (anonymous):

\[r'(t) = -asin(t)i+acos(t)j\]

OpenStudy (jamesj):

yep

OpenStudy (anonymous):

so then after calculating ||r'(t)|| I end up with just a, want me to show that work?

OpenStudy (jamesj):

no. that's right: ||r'|| = a

OpenStudy (jamesj):

Hence T = r' / ||r'|| = ...?

OpenStudy (anonymous):

-sin(t)i+cos(t)j

OpenStudy (jamesj):

yes

OpenStudy (jamesj):

hence T' = ....

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

1/a

OpenStudy (anonymous):

You're a genious

OpenStudy (anonymous):

T' doesn't = 1/a

OpenStudy (anonymous):

But I see how (||T' (t) ||)/(||r' (t) ||) will eventually = 1/a now

OpenStudy (jamesj):

T' = -cos t i - sin t j ==> ||T'|| = 1. Therefore k = ||T'||/||r'|| = 1/a

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

Thanks so much

OpenStudy (anonymous):

Are you a math major?

OpenStudy (jamesj):

I was, yes.